A bullet is fired normallyon a wodden plank,which is immovable.it losses 25% of it's momebtum in penetrating a thickness of 3.5cm.the total thickness penetrated by bullet is-

Grade:

2 Answers

piss on you
11 years ago

Mass is constant. So, 25% loss in momentum => 25% loss in velocity

Initial velocity u becomes 3u/4

                                             2aS=v2-u2

S = 3.5cm, u=u, v=3u/4

7a = (9u2/16) - u2

i.e 7a = -7u2/16  or a = -u2/16

After coming to rest, v = 0 , u = u , a = -u2/16

2aD = 0 - u2

2(-u2/16)D = -u2

i.e.    D = 8 cm

neeraj agarwal
11 years ago

energy hisab kitab                                        1/16  of energy works=3.5cm                                                                                                                                       9/16 of energy will work for 9*3.5=31.5cm                                                                                                                                        net distance=31.5+3.5=35cm